PK2BM(1) PK2BM(1) NAME pk2bm−createabitmapfromaTeXpkfont SYNOPSIS: pk2bm[-bh][-Wwidth][-Hheight]-ccharpkfont pk2bm[-bh][-Wwidth][-Hheight]-ooctcharpkfont DESCRIPTION Thisprogramgeneratesabitmap(ASCIIfile)whichcanbeusedtocreateX11applications.The bitmapfileconsistsofallpixelsofthespecifiedcharacter(viathe-cor-ooption)fromthegiven pkfont.Theformatisdescribedinbitmap(X11). Thepkfontisapackedfontfilegeneratedbygftopk(TeX)fromagffont.Agffontistheoutputof METAFONTaprogramtodesignfontsinadeviceindependentway. Withthe-bflagabitmapisgeneratedinwhichallblackpixelsaredrawnasa‘*’andallwhite bitsasa‘.’.Withthe-hflagahexadecimalbitmapdumpisgenerated. The-Wand/or-Hoptionscanbeusedtocreateabitmapofrespectivally‘width’and‘height’ pixels.Thepk-bitmapwillinthiscasebecenteredaccordingtothesenewdimensions. Theoutputiswrittentothestandardoutput. SEEALSO ‘METAFONT’,DonaldKnuth. ‘ThePKtypeprocessor’,belongingtotheMETAFONTware. bitmap(X11), gftopk(TeX), pktype(TeX), ps2pk(1) AUTHOR PietTutelaers rcpt@urc.tue.nl TeX 1